# Measuring the Measures: Discriminative Capacity of Representational Similarity Metrics Across Model Families

### v3

@@ 1 Introduction
Representational similarity metrics have become fundamental tools for understanding deep neural networks, enabling comparisons across architectures, layers, and even between artificial and biological vision systems. The field has developed diverse approaches to quantify representational similarity, including methods that utilize stimulus-by-stimulus similarity matrices to compare across networks, like Representational Similarity Analysis (RSA) as well as methods that learn explicit mappings to align neural dimensions, such as linear predictivity, Procrustes alignment proc1 ; proc2 , and soft matching softmatching . Whie these metrics have advanced our understanding of representational alignment, a critical question remains unexplored: how well do these metrics discriminate between different model families?
@@ 4 Discussion
Our findings highlight that representational similarity metrics differ in their ability to seperate model families. Metrics imposing stronger alignment constraints provide higher discriminability than looser measures, and non-fitting approaches show strong separation. These findings provide practical guidance for representation analysis: researchers should select metrics based on their discrimination goals rather than defaulting to popular choices. More broadly, our framework offers a principled way to benchmark similarity metrics and clarify their trade-offs, paving the way for more interpretable and goal-aligned comparisons across models and between models and brains.
@@ Appendix B Model Family and Architecture Choices
Label-free training with Transformer backbones. We include DINO-ViT-B/16 and DINO-ViT-S/16 [22], MoCo-ViT-B/16 [21], and MAE-ViT-B/16 and MAE-ViT-L/16 [25]. “‘

### v4

@@ 1 Introduction
Representational similarity metrics have become fundamental tools for understanding deep neural networks, enabling comparisons across architectures, layers, and even between artificial and biological vision systems. The field has developed diverse approaches to quantify representational similarity, including methods that utilize stimulus-by-stimulus similarity matrices to compare across networks, like Representational Similarity Analysis (RSA) as well as methods that learn explicit mappings to align neural dimensions, such as linear predictivity, Procrustes alignment proc1 ; proc2 , and soft matching softmatching . While these metrics have advanced our understanding of representational alignment, a critical question remains unexplored: how well do these metrics discriminate between different model families?
@@ 4 Discussion
Our findings highlight that representational similarity metrics differ in their ability to separate model families. Metrics imposing stronger alignment constraints provide higher discriminability than looser measures, and non-fitting approaches show strong separation. One limitation of our analysis is that we are using human intuitions to impose a desideratum for evaluating metrics: that models within the same family should cluster together and separately from models from another family. However, sometimes representational convergence may emerge even between models from different families. Nevertheless, our findings still provide practical guidance for representation analysis: researchers should select metrics based on their discrimination goals rather than defaulting to popular choices. More broadly, our framework offers a principled way to benchmark similarity metrics and clarify their trade-offs, paving the way for more interpretable and goal-aligned comparisons across models and between models and brains.
@@ Appendix B Model Family and Architecture Choices
Label-free training with Transformer backbones. We include DINO-ViT-B/16 and DINO-ViT-S/16 [22], MoCo-ViT-B/16 [21], and MAE-ViT-B/16 and MAE-ViT-L/16 [25].